### v3

Abstract
We introduce a new global sensitivity measure, the global activity scores. We establish its theoretical connection with Sobol’ sensitivity indices and demonstrate its performance through numerical examples. In these examples, we compare global activity scores with Sobol’ sensitivity indices, derivative-based sensitivity measures, and activity scores. The results show that in the presence of noise or high variability, global activity scores outperform derivative-based measures and activity scores, while in noiseless settings the three approaches yield similar results.
keywords:
Global sensitivity analysis, Sobol’ sensitivity indices, derivative-based global sensitivity measures, activity scores
{AMS}
65C05, 65C20
1 Introduction
Global sensitivity analysis quantifies how uncertainty in a model’s output can be attributed to uncertainties in its inputs. The insights gained from such an analysis make it possible to identify the most influential input parameters, enabling the modeler, among other benefits, to reduce model complexity by fixing variables that are not deemed important. Applications of global sensitivity analysis are widespread, spanning the natural sciences, engineering, social sciences, and mathematical sciences. Several approaches have been proposed in the literature, including Sobol’ sensitivity indices, derivative-based global sensitivity measures, and activity scores. In this paper, we introduce a new sensitivity measure, the global activity scores. We investigate its theoretical relationship with Sobol’ indices and present numerical results illustrating its advantages over existing measures, particularly in the presence of noise.
The remainder of the paper is organized as follows. Section 2 reviews Sobol’ sensitivity indices (Sobol’ [20]), derivative-based global sensitivity measures (Sobol’ and Kucherenko [19]), and activity scores (Constantine et al. [4]). Section 3 introduces our new sensitivity measure, global activity scores. Section 4 presents numerical experiments applying these methods to selected examples. We conclude in Section 5.
2 Global Sensitivity Measures
In this section we review some popular types of sensitivity measures, in particular, the variance-based and derivative-based measures. For a comprehensive survey of sensitivity measures, see Iooss and Lemaître [8].
2.1 Sobol’ Sensitivity Indices
Consider a $d$-dimensional input vector $\boldsymbol{x}=(x_{1},\ldots,x_{d})$, a square-integrable function $f(\boldsymbol{x})$ defined on $(0,1)^{d}$, and let the index set be defined as $D=\{1,2,\ldots,d\}$. The ANOVA decomposition of $f(\boldsymbol{x})$ is
$f(\boldsymbol{x})=\sum_{u\subseteq D}f_{u}(\boldsymbol{x}^{u}),$
where $f_{u}(\boldsymbol{x}^{u})$ is the component function that only depends on $\boldsymbol{x}^{u}$. For the empty set, we put $f_{\emptyset}=\int f(\boldsymbol{x})d\boldsymbol{x}$.
If we assume $\boldsymbol{x}$ has uniform distribution on $(0,1)^{d}$, we can write
$\mathbb{E}[f(\boldsymbol{x})]=\int_{(0,1)^{d}}f(\boldsymbol{x})d\boldsymbol{x},$
and
$\text{Var}(f(\boldsymbol{x}))=\sigma^{2}=\int_{(0,1)^{d}}f^{2}(\boldsymbol{x})d\boldsymbol{x}-\mathbb{E}[f(\boldsymbol{x})]^{2}.$
As a consequence of the orthogonality of the ANOVA decomposition, the variance of $f$ can be written as
$\sigma^{2}=\sum_{u\subseteq D}\sigma_{u}^{2},$
where $\sigma_{u}^{2}$ is the variance for the component function $f_{u}$:
$\sigma_{u}^{2}=\int_{(0,1)^{d}}f_{u}^{2}(\boldsymbol{x})d\boldsymbol{x}-\left(\int_{(0,1)^{d}}f_{u}(\boldsymbol{x})d\boldsymbol{x}\right)^{2}=\int_{(0,1)^{d}}f_{u}^{2}(\boldsymbol{x})d\boldsymbol{x}.$
The Sobol’ sensitivity indices for the subset $u$ are defined as
[…]
where $\underline{S}_{u}$ is called the lower Sobol’ sensitivity index (or, the main effect) and $\overline{S}_{u}$ is called the upper Sobol’ sensitivity index (or, the total effect). If $\underline{S}_{u}$ is close to 1, then the parameters $\boldsymbol{x}^{u}$ are viewed as important to the model output. In practice the case when $u$ is a singleton $\{i\}$ is often considered. If $\overline{S}_{\{i\}}$ is close to 0, we consider $x_{i}$ to be an unimportant variable, and freeze it to its mean value to simplify the model $f$.
Upper Sobol’ indices can be written as (Sobol’ [20])
(1) $\bar{S}_{i}=\frac{1}{2\sigma^{2}}\int{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{z}dv_{i}},$
where $(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{-\{i\}})$ means the vector whose $i$th component is $v_{i}$, and whose $j$th component $(j\neq i)$ is $z_{j}$. Eqn. (1) can be generalized to random vectors $\boldsymbol{x}=(x_{1},\ldots,x_{d})$ with distribution function $\boldsymbol{F}(\boldsymbol{x})=F_{1}(x_{1})\cdots F_{d}(x_{d})$ (Kucherenko et al. [12]) as
(2) $\bar{S}_{i}=\frac{1}{2\sigma^{2}}\int{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{F}(\boldsymbol{z})dF_{i}(v_{i})}.$
2.2 Derivative-based sensitivity measures
Computationally more efficient sensitivity measures can be obtained if the partial derivatives of $f$ exist. For example, Campolongo et al. [2] introduced a global sensitivity measure based on $\int_{(0,1)^{d}}\left|\frac{\partial f(\boldsymbol{x})}{\partial x_{i}}\right|d\boldsymbol{x}$, and Sobol’ and Kucherenko [19] introduced a global sensitivity measure based on $\int_{(0,1)^{d}}(\frac{\partial f(\boldsymbol{x})}{\partial x_{i}})^{2}d\boldsymbol{x}$. Here we discuss the second measure further.
The derivative-based global sensitivity measure (DGSM) of Sobol’ and Kucherenko [19] is given by
(3) $v_{i}=\int_{(0,1)^{d}}\left(\frac{\partial f(\boldsymbol{x})}{\partial x_{i}}\right)^{2}d\boldsymbol{x}=\mathbb{E}\left[\left(\frac{\partial f(\boldsymbol{x})}{\partial x_{i}}\right)^{2}\right].$
Sobol’ [18] showed that if $f(\boldsymbol{x})$ is a linear function on each of its components $x_{i}$, then
(4) $\overline{S}_{i}=\frac{1}{12}\frac{v_{i}}{\sigma^{2}}.$
In general, Sobol’ and Kucherenko [19] established the following inequality
(5) $\overline{S}_{i}\leq\frac{1}{\pi^{2}}\frac{v_{i}}{\sigma^{2}}.$
DGSM can be generalized to nonuniform distributions, where the random vector $\boldsymbol{x}=(x_{1},\ldots x_{d})$ has independent components with marginal distributions $F_{1},\ldots,F_{d}$, and each $F_{i}$ has a corresponding density function $f_{i}$. Sobol’ and Kucherenko [19] and Kucherenko and Iooss [11] obtained several results in this setting. In particular, the following generalization of inequality (5) was obtained in [11] for distributions that satisfy certain regularity conditions:
(6) $\overline{S}_{i}\leq D(F_{i})\frac{v_{i}}{\sigma^{2}},$
where
$D(F_{i})=4\left[\sup_{x\in\mathbb{R}}\frac{\min(F_{i}(x),1-F_{i}(x))}{f_{i}(x)}\right]^{2}.$
2.3 Activity scores
The active subspace (AS) method (Constantine et al. [4]) finds the important directions of a function using the gradient of the function, and uses the important directions to reduce the domain of the function to a subspace. The activity score, introduced by Constantine and Diaz [3], is a global sensitivity measure obtained from this information.
Consider a square-integrable function $f(\boldsymbol{x})$ defined on $(0,1)^{d}$, with finite partial derivatives that are square-integrable. Let $p(\boldsymbol{x})$ be the uniform probability density function on $(0,1)^{d}$ and consider the gradient of $f$
$\nabla f(\boldsymbol{x})=\left[\frac{\partial f(\boldsymbol{x})}{\partial x_{1}},\frac{\partial f(\boldsymbol{x})}{\partial x_{2}},\ldots,\frac{\partial f(\boldsymbol{x})}{\partial x_{d}}\right]^{T}.$
Define the matrix $C_{\text{as}}$ as
(7) $C_{\text{as}}=\mathbb{E}[\nabla f(\boldsymbol{x})\nabla f(\boldsymbol{x})^{T}],$
where the expectation is computed using the density function $p$. Let the eigenvalue decomposition of $C_{\text{as}}$ be
$C_{\text{as}}=\textbf{W}\Lambda_{\text{as}}\textbf{W}^{T},$
where $\textbf{W}=[\textbf{w}_{1},\ldots,\textbf{w}_{d}]$ is the $d\times d$ orthogonal matrix of eigenvectors, and $\Lambda_{\text{as}}=\text{diag}(\lambda_{1},\ldots,\lambda_{d})$ with $\lambda_{1}\geq\ldots\geq\lambda_{d}\geq 0$ is the diagonal matrix of eigenvalues in descending order.
Matrix $C_{\text{as}}$ can be approximated using the Monte Carlo method as
(8) $C_{\text{as}}\approx\hat{C}_{\text{as}}=\frac{1}{N}\sum_{j=1}^{N}(\nabla_{\boldsymbol{x}}f(\boldsymbol{x}^{({j})}))(\nabla_{\boldsymbol{x}}(f(\boldsymbol{x}^{(j)})))^{T},$
where the $\boldsymbol{x}^{(1)},\ldots,\boldsymbol{x}^{(N)}$ are a random sample of size $N$ from the density $p$.
If there exists an integer $m$ such that eigenvalues $\lambda_{m+1},\ldots,\lambda_{d}$ are sufficiently small, then the active subspace method approximates $f(\boldsymbol{x})$ with a lower dimensional function $g(\textbf{W}_{1}^{T}\textbf{x})$ where $\textbf{W}_{1}$ is the $d\times m$ matrix containing the first $m$ eigenvectors. The dimension of $g$ is $m$, whereas the dimension of $f$ is $d$.
The activity score for the $i$th parameter is defined as
(9) $\alpha_{i}(m)=\sum_{j=1}^{m}\lambda_{j}w_{ij}^{2},$
where $\textbf{w}_{j}=[w_{1j},\ldots,w_{dj}]^{T}$ is the $j$th eigenvector, $m\leq d$, and $i=1,\ldots,d$.
Constantine and Diaz [3] showed that the activity scores are bounded by DGSM
$\alpha_{i}(m)\leq v_{i},$
where $i=1,\ldots,d$, and the inequality becomes an equality if $m=d$.
The following inequality between the Sobol’ sensitivity index and activity scores was also established in [3]:
$\overline{S}_{i}\leq\frac{1}{\pi^{2}}\frac{\alpha_{i}(m)+\lambda_{m+1}}{\sigma^{2}},$
where $i=1,\ldots,d$, $m=1,\ldots,d-1$. Duan and Ökten [5] note that this result can be generalized to non-uniform distributions over $\mathbb{R}^{d}$ using Eqn. (6) as
$\overline{S}_{i}\leq D(F_{i})\frac{\alpha_{i}(m)+\lambda_{m+1}}{\sigma^{2}},$
where $F_{i}$ and $f_{i}$, $i=1,\ldots,d$, are marginal distributions and density functions for the components of the random input vector $\boldsymbol{x}$.
3 Global Activity Scores
In this section, we introduce a new global sensitivity measure: global activity scores. This measure replaces the expectations of gradients used in traditional activity scores with expectations of finite differences. We will demonstrate, through numerical experiments on noisy problems, the advantages of using finite differences over gradients.
Consider a square-integrable real-valued bounded function $f(\boldsymbol{z})$ with domain $\Omega\subset\mathbb{R}^{d}$ and continuous second-order partial derivatives. Suppose $\Omega$ is endowed with a probability measure with a cumulative distribution function in the form $\boldsymbol{F}(\boldsymbol{z})=F_{1}(z_{1})\cdot\ldots\cdot F_{d}(z_{d})$, where $F_{i}$ are marginal distribution functions.
Let $D_{\boldsymbol{z}}$ be an operator, acting on $f$, whose value when evaluated at $(\boldsymbol{v},\boldsymbol{z})$ is given by:
$D_{\boldsymbol{z}}f(\boldsymbol{v},\boldsymbol{z})=[D_{\boldsymbol{z},1}f(v_{1},\boldsymbol{z}),\ldots,D_{\boldsymbol{z},d}f(v_{d},\boldsymbol{z})]^{T},$
where
(10) $D_{\boldsymbol{z},i}f(v_{i},\boldsymbol{z})=(f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{-\{i\}})-f(\boldsymbol{z}))/(v_{i}-z_{i}).$
The notation $(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{-\{i\}})$ means the vector whose $i$th component is $v_{i}$, and whose $j$th component $(j\neq i)$ is $z_{j}$. The notations $\boldsymbol{v}_{\{i\}}$ and $v_{i}$ both mean the $i$th component of $\boldsymbol{v}$. We use the former notation only when we are splicing and replacing a component from a vector with another.
Define the $d\times d$ matrix $\boldsymbol{C}_{\text{gas}}$ by
(11) $\boldsymbol{C}_{\text{gas}}={\rm E}\,[{\rm E}\,[(D_{\boldsymbol{z}}f)(D_{\boldsymbol{z}}f)^{T}|\boldsymbol{z}]],$
where the inner conditional expectation fixes $\boldsymbol{z}$ and averages over the components of $\boldsymbol{v}=[v_{1},\ldots,v_{d}]^{T}$, and the outside expectation averages with respect to $\boldsymbol{z}$. Both $\boldsymbol{v}$ and $\boldsymbol{z}$ follow the same continuous probability distribution $\boldsymbol{F}$.
We next discuss the necessary conditions for the existence of the integrals in $\boldsymbol{C}_{\text{gas}}$, for the case of bivariate $f$ - the proofs generalize to higher dimensions in a straightforward way. Let $\tilde{F}(x)=\prod_{i\in S}F_{i}(x_{i})$ where $S$ is a list of indices $i_{1},\ldots,i_{s}$, $s\leq d$, and each $i_{k}$, $k=1,\ldots,s$, belongs to the set of indices $\{1,\ldots,d\}$. Then $\tilde{F}$ is an $s$-dimensional Stieltjes measure function (Kingman and Taylor [9]). The corresponding measure $\mu_{\tilde{F}}$ takes the following value on $s$-dimensional intervals $a,b)=[a_{1},b_{1})\times\ldots\times[a_{s},b_{s})$:
$\mu_{\tilde{F}}([a,b))=\prod_{k=1,\ldots,s}(F_{i_{k}}(b_{i_{k}})-F_{i_{k}}(a_{i_{k}})).$
Using standard measure theory techniques $\mu_{\tilde{F}}$ can be extended to Borel sets in $\mathbb{R}^{s}$, whose completion is the Lebesgue-Stieltjes (LS) measurable sets in $\mathbb{R}^{s}$.
Theorem 3.1.
Let $f$ be a bounded function with continuous first-order partial derivatives. Define $g(v_{1},z_{1},z_{2})$ by
$g(v_{1},z_{1},z_{2})=\begin{cases}\left(\frac{f(v_{1},z_{2})-f(z_{1},z_{2})}{v_{1}-z_{1}}\right)^{2}\text{if }v_{1}\neq z_{1},\\ \left(\frac{\partial f}{\partial z_{1}}(z_{1},z_{2})\right)^{2}\text{if }v_{1}=z_{1}.\end{cases}$
Then the integral $C_{11}$, defined as the limit of Riemann-Stieltjes integrals,
[…]
exists, where
$\mathcal{B}_{n}=\{(v_{1},z_{1},z_{2})\in\mathbb{R}^{3},\text{s.t. }|v_{1}-z_{1}|<1/n\},n>0.$
Proof 3.2.
For each $n>0$, the function $g$ is bounded and continuous on $\mathbb{R}^{3}\setminus\mathcal{B}_{n}$, hence the Riemann–Stieltjes integral
$I_{n}=\int_{\mathbb{R}^{3}\setminus\mathcal{B}_{n}}g(v_{1},z_{1},z_{2})\,dF_{1}(v_{1})\,dF_{1}(z_{1})\,dF_{2}(z_{2})$
is well-defined. The sequence $(I_{n})$ is monotone increasing, since $\mathcal{B}_{n+1}\subseteq\mathcal{B}_{n}$, and it is bounded above by
$\int_{\mathbb{R}^{3}}g\,d\tilde{F},\qquad\tilde{F}=F_{1}\times F_{1}\times F_{2}.$
Thus the limit $\lim_{n\to\infty}I_{n}$ exists. Moreover, since $g$ is bounded on $\mathbb{R}^{3}$, say $0\leq g\leq C$, and $F_{1},F_{2}$ are continuous, we have
$\int_{\mathcal{B}_{n}}g\,d\tilde{F}\leq C\,\mu_{\tilde{F}}(\mathcal{B}_{n})\;\longrightarrow\;0,$
as $n\to\infty$. Therefore,
$C_{11}=\int_{\mathbb{R}^{3}}g\,d\tilde{F}$.
Theorem 3.3.
Let $f$ be a bounded function with continuous first-order partial derivatives. Define
$g(v_{1},v_{2},z_{1},z_{2})=\frac{(f(v_{1},z_{2})-f(z_{1},z_{2}))(f(z_{1},v_{2})-f(z_{1},z_{2}))}{(v_{1}-z_{1})(v_{2}-z_{2})},$
if $v_{1}\neq z_{1}$ and $v_{2}\neq z_{2}$. If $v_{1}=z_{1}$, then the corresponding term $\frac{f(v_{1},z_{2})-f(z_{1},z_{2})}{v_{1}-z_{1}}$ is replaced by $\frac{\partial f}{\partial z_{1}}(z_{1},z_{2})$. The case $v_{2}=z_{2}$ is handled similarly. Then the integral $C_{12}$, defined as the limit of Riemann–Stieltjes integrals,
[…]
exists, where, for $n>0$,
$\mathcal{B}_{n}=\{(v_{1},v_{2},z_{1},z_{2})\in\mathbb{R}^{4}:|v_{1}-z_{1}|<1/n\}\,\cup\,\{(v_{1},v_{2},z_{1},z_{2})\in\mathbb{R}^{4}:|v_{2}-z_{2}|<1/n\}.$
Proof 3.4.
[…]
and
[…]
where $\tilde{F}=F_{1}\times F_{2}\times F_{1}\times F_{2}.$ Then $(J_{n})$ is an increasing sequence bounded by $\int_{\mathbb{R}^{4}}|g|\,d\tilde{F}$, so $\lim_{n\to\infty}J_{n}$ exists. Moreover,
$0\leq|g(v_{1},v_{2},z_{1},z_{2})|-g(v_{1},v_{2},z_{1},z_{2})\leq 2|g(v_{1},v_{2},z_{1},z_{2})|,$
which implies
[…]
exists, since the sequence is increasing and bounded by $2\int_{\mathbb{R}^{4}}|g|d\tilde{F}$. Since the limits of the integrals of $|g|$ and |g|$g|$-g both exist, the limit
$C_{12}=\lim_{n\rightarrow\infty}I_{n}$
exists as well. Finally, since $g$ is bounded and $F_{i}$ are continuous ($i=1,2$), $\int_{\mathcal{B}_{n}}gd\tilde{F}\rightarrow 0$ as $n\rightarrow\infty$, and hence
$C_{12}=\int_{\mathbb{R}^{4}}g(v_{1},v_{2},z_{1},z_{2})\,d\tilde{F}$.
Next we discuss the estimation of $\boldsymbol{C}_{\text{gas}}$. Each entry of the matrix $\boldsymbol{C}_{\text{gas}}$ is a double expectation. The inner expectations on the diagonal of $\boldsymbol{C}_{\text{gas}}$ can be expressed as a one-dimensional integral, which we approximate using a deterministic quadrature rule. The outer expectation is a multidimensional integral, estimated by Monte Carlo. For example, consider a bivariate function and the first entry of the matrix:
$C_{11}$ $=\int_{\mathbb{R}^{3}}\left(\frac{f(v_{1},z_{2})-f(z_{1},z_{2})}{v_{1}-z_{1}}\right)^{2}dF_{1}(v_{1})d\boldsymbol{F}(\boldsymbol{z}).$
To estimate this integral, generate a sample $\boldsymbol{z}^{(i)}=(z_{1}^{(i)},z_{2}^{(i)})$ from $\boldsymbol{F}$, for $i=1,\ldots,N.$ The corresponding inner integral is
$\int_{\mathbb{R}}\left(\frac{f(v_{1},z^{(i)}_{2})-f(z^{(i)}_{1},z^{(i)}_{2})}{v_{1}-z^{(i)}_{1}}\right)^{2}dF_{1}(v_{1}).$
This integral has a singularity at $v_{1}=z^{(i)}_{1}$. To handle it, split the domain at the singularity into $(-\infty,z^{(i)}_{1})$ $(z^{(i)}_{1},\infty), and$ approximate each part using a deterministic quadrature rule such as Newton-Cotes or Gauss-Legendre. Repeating this procedure for $i=1,\ldots,N$ and averaging the results yields an estimate of $C_{11}$.
Consider the eigenvalue decomposition of $\boldsymbol{C}_{\text{gas}}$
(12) $\boldsymbol{C}_{\text{gas}}=\boldsymbol{U}\Lambda_{\text{gas}}\boldsymbol{U}^{T},$
where $\boldsymbol{U}=[\textbf{u}_{1},\ldots,\textbf{u}_{d}]$ is the $d\times d$ orthogonal matrix of eigenvectors, and $\Lambda_{\text{gas}}=\text{diag}(\lambda_{1},\ldots,\lambda_{d})$ with $\lambda_{1}\geq\ldots\geq\lambda_{d}\geq 0$ is the diagonal matrix of eigenvalues in descending order. The global activity scores are defined based on the eigenvalue decomposition of $\boldsymbol{C}_{\text{gas}}$.
Definition 3.5.
The global activity score for the $i$th parameter, $1\leq i\leq d$, is
(13) $\gamma_{i}(m)=\sum_{j=1}^{m}{\lambda_{j}u_{ij}^{2}},$
where $\lambda_{1},\ldots,\lambda_{d}$ are the eigenvalues from $\Lambda_{\text{gas}}$, $\boldsymbol{u}_{j}=[u_{1j},\ldots,u_{dj}]^{T}$ is the $j$th eigenvector from $\boldsymbol{U}$, and $m\leq d$.
The next theorem presents an inequality between the upper Sobol’ sensitivity index $\bar{S}_{i}$ and $\gamma_{i}(d)$.
Theorem 3.6.
If $\Omega=(0,1)^{d}$, endowed with the uniform probability measure, then
(14) $\bar{S}_{i}\leq\frac{1}{2}\frac{\gamma_{i}(d)}{\sigma^{2}},$
for $i=1,\ldots,d$.
Proof 3.7.
Consider the upper Sobol’ index $\bar{S}_{i}$ (Eqn. (1))
$\bar{S}_{i}=\frac{1}{2\sigma^{2}}\int{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{z}dv_{i}}.$
Since $(v_{i}-z_{i})^{2}\leq 1$ for any $v_{i},z_{i}$ in $(0,1)$, we have
$\bar{S}_{i}\leq\frac{1}{2\sigma^{2}}\int{((f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))/(z_{i}-v_{i}))^{2}}d\boldsymbol{z}dv_{i}.$
Let $\boldsymbol{\bar{S}}=[\bar{S}_{1},...,\bar{S}_{d}]^{T}$. We have
$\boldsymbol{\bar{S}}\leq\frac{1}{2\sigma^{2}}diag(\boldsymbol{C}_{\text{gas}})=\frac{1}{2\sigma^{2}}diag(\boldsymbol{U}\Lambda_{\text{gas}}\boldsymbol{U}^{T})$
where $diag(A)$ is the vector obtained from the diagonal elements of the matrix $A$. Since the $i$th value of $diag(\boldsymbol{U}\Lambda_{\text{gas}}\boldsymbol{U}^{T})$ is $\gamma_{i}(d)$ by definition of $\gamma_{i}(d)$, the above equality implies
$\bar{S}_{i}\leq\frac{1}{2\sigma^{2}}\gamma_{i}(d)$
for $i=1,\ldots,d.$
The following result links any $\gamma_{i}(m),1\leq m\leq d-1$, to the upper Sobol’ index.
Corollary 3.8.
If $\Omega=(0,1)^{d}$, endowed with the uniform probability measure, then
(15) $\bar{S}_{i}\leq\frac{1}{2}\frac{\gamma_{i}(m)+\lambda_{m+1}}{\sigma^{2}},$
where $i=1,\ldots,d$, and $\lambda_{m+1}$ is the $(m+1)$th eigenvalue of matrix $\boldsymbol{C}_{\text{gas}}$.
Proof 3.9.
The proof follows from Theorem 3.6, the orthogonality of $\boldsymbol{U}$, and the following inequality:
$\gamma_{i}(d)=\gamma_{i}(m)+\sum_{j=m+1}^{d}{\lambda_{j}u_{ij}^{2}}\leq\gamma_{i}(m)+\lambda_{m+1}\sum_{j=m+1}^{d}{u_{ij}^{2}}.$
Theorem 3.6 and Corollary 3.8 can be generalized to unbounded domains and nonuniform measures. Let $\Omega=\mathbb{R}^{d}$, endowed with a probability distribution function in the form $\boldsymbol{F}(\boldsymbol{z})=F_{1}(z_{1})\cdot\ldots\cdot F_{d}(z_{d})$.
Theorem 3.10.
Let $f(\boldsymbol{z})$ be a bounded function on $\Omega$. For any $0<\epsilon<1$
(16) $\bar{S}_{i}\leq\frac{(b^{\prime}-a^{\prime})^{2}}{2}\frac{\gamma_{i}(d)+\kappa}{\sigma^{2}},$
where $a^{\prime},b^{\prime}\in\mathbb{R}$ are such that $\int_{(a^{\prime},b^{\prime})^{d}}d\boldsymbol{F}(\boldsymbol{z})=1-\epsilon$, $i=1,\ldots,d$, and $\kappa$ is a positive constant given by
(17) $\kappa=\frac{2\epsilon-\epsilon^{2}}{(b^{\prime}-a^{\prime})^{2}}\sup_{\boldsymbol{z},\boldsymbol{v}\in\Omega}(f(\boldsymbol{z})-f(\boldsymbol{v}))^{2}.$
Proof 3.11.
From the definition of generalized upper Sobol’ indices (Eqn. (2)), we have
$\bar{S}_{i}$ $=\frac{1}{2\sigma^{2}}\int{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{F}(\boldsymbol{z})dF_{i}(v_{i})}$
$=\frac{1}{2\sigma^{2}}\int_{\Omega\times\Omega}{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{F}(\boldsymbol{z})d\boldsymbol{F}(\boldsymbol{v})}.$
Let $\Omega^{\prime}=(a^{\prime},b^{\prime})^{d}$, and write $\Omega\times\Omega$ as the union of $\Omega^{\prime}\times\Omega^{\prime}$ and the complement $(\Omega^{\prime}\times\Omega^{\prime})^{c}$. Since $(z_{i}-v_{i})^{2}\leq(b^{\prime}-a^{\prime})^{2}$, for any $z_{i},v_{i}\in(a^{\prime},b^{\prime})$, we have
$\int_{\Omega^{\prime}\times\Omega^{\prime}}{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{F}(\boldsymbol{z})d\boldsymbol{F}(\boldsymbol{v})}$
$\leq(b^{\prime}-a^{\prime})^{2}\int_{\Omega^{\prime}\times\Omega^{\prime}}{((f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{-\{i\}}))/(z_{i}-v_{i}))^{2}d\boldsymbol{F}(\boldsymbol{z})d\boldsymbol{F}(\boldsymbol{v})}.$
The integral over the complement of $\Omega^{\prime}\times\Omega^{\prime}$ can be bounded by
$\int_{(\Omega^{\prime}\times\Omega^{\prime})^{c}}{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{F}(\boldsymbol{z})d\boldsymbol{F}(\boldsymbol{v})}\leq(b^{\prime}-a^{\prime})^{2}\kappa.$
Using the bounds for the two integrals and dividing by $2\sigma^{2}$ we obtain
(18) $\bar{S}_{i}$ $\leq\frac{(b^{\prime}-a^{\prime})^{2}}{2\sigma^{2}}\int_{\Omega\times\Omega}{((f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}:\boldsymbol{z}_{-\{i\}}))/(z_{i}-v_{i}))^{2}d\boldsymbol{F}(\boldsymbol{z})d\boldsymbol{F}(\boldsymbol{v})}+\frac{(b^{\prime}-a^{\prime})^{2}}{2\sigma^{2}}\kappa$
$=\frac{(b^{\prime}-a^{\prime})^{2}}{2\sigma^{2}}\int_{\Omega\times\mathbb{R}}{((f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}:\boldsymbol{z}_{-\{i\}}))/(z_{i}-v_{i}))^{2}d\boldsymbol{F}(\boldsymbol{z})dF_{i}(v_{i})}+\frac{(b^{\prime}-a^{\prime})^{2}}{2\sigma^{2}}\kappa.$
Let $\boldsymbol{\bar{S}}=[\bar{S}_{1},\ldots,\bar{S}_{d}]^{T}$. Inequality (18) can be written as
$\boldsymbol{\bar{S}}$ $\leq\frac{(b^{\prime}-a^{\prime})^{2}}{2\sigma^{2}}\left(diag(\boldsymbol{C}_{\text{gas}})+\kappa\boldsymbol{1_{d}}\right)=\frac{(b^{\prime}-a^{\prime})^{2}}{2\sigma^{2}}\left(diag(\boldsymbol{U}\Lambda_{\text{gas}}\boldsymbol{U}^{T})+\kappa\boldsymbol{1_{d}}\right),$
where $\boldsymbol{1_{d}}$ denotes the $d$-dimensional vector of ones. Since $\gamma_{i}(d)$ is the $i$th value of $diag(\boldsymbol{U}\Lambda_{\text{gas}}\boldsymbol{U}^{T})$, the proof is over.
Remark 3.12.
If the domain of $f$ is a finite rectangle $(a,b)^{d}$, then the integral over the unbounded domain in the proof of Theorem 3.10 vanishes, and we obtain the bound
$\bar{S}_{i}\leq\frac{(b-a)^{2}}{2}\frac{\gamma_{i}(d)}{\sigma^{2}}.$
If we put $a=0,b=1$ in the above inequality, we obtain the bound in Theorem 3.6.
For any $m$, $1\leq m\leq d-1$, we have the following inequality for $\gamma_{i}(m)$ and $\bar{S}_{i}$.
Corollary 3.13.
With the same assumptions as in Theorem 3.10, we have
(19) $\bar{S}_{i}\leq\frac{(b^{\prime}-a^{\prime})^{2}}{2}\frac{\gamma_{i}(m)+\lambda_{m+1}+\kappa}{\sigma^{2}},$
where $i=1,\ldots,d$ and $\lambda_{m+1}$ is the $(m+1)$th eigenvalue of matrix $\boldsymbol{C}_{\text{gas}}$.
Proof 3.14.
The proof follows from Theorem 3.10, as in the proof of Corollary 3.8.
Remark 3.15.
If the domain of $f$ is a finite rectangle $(a,b)^{d}$, then the bound in Corollary 3.13 simplifies as
$\bar{S}_{i}\leq\frac{(b-a)^{2}}{2}\frac{\gamma_{i}(m)+\lambda_{m+1}}{\sigma^{2}}.$
If we put $a=0,b=1$ in the above inequality, we obtain the bound in Corollary 3.8.
The next theorem presents a much simpler relationship between upper Sobol’ indices and global activity scores for quadratic functions.
Theorem 3.16.
If $\Omega=\mathbb{R}^{d}$ and $\boldsymbol{z}$ has the multivariate standard normal probability distribution $\boldsymbol{F}=(F_{1},\ldots,F_{d})$, and $f(\boldsymbol{z})=\frac{1}{2}\boldsymbol{z}^{T}\boldsymbol{Az}+\boldsymbol{b}^{T}\boldsymbol{z}$, where $\boldsymbol{A}$ is a symmetric $d\times d$ matrix, then
(20) $\bar{S}_{i}=\frac{\gamma_{i}(d)}{\sigma^{2}},$
for $i=1,\ldots,d$.
Proof 3.17.
We have $\boldsymbol{z}=(z_{1},\ldots,z_{d})^{T}$, and $\boldsymbol{b}=(b_{1},\ldots,b_{d})$. Then
$2\sigma^{2}\bar{S}_{i}$ $=\int{(f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))^{2}d\boldsymbol{F}(\boldsymbol{z})dF_{i}(v_{i})}$
$=\int{\left(\frac{\partial f(\boldsymbol{x}_{i})}{\partial z_{i}}\right)^{2}(z_{i}-v_{i})^{2}d\boldsymbol{F}(\boldsymbol{z})dF_{i}(v_{i})}$
$={\rm E}\,\left[\left(\frac{\partial f(\boldsymbol{x}_{i})}{\partial z_{i}}\right)^{2}(z_{i}-v_{i})^{2}\right],$
where $\boldsymbol{x}_{i}=(z_{1},\ldots,z_{i-1},\frac{z_{i}+v_{i}}{2},z_{i+1},\ldots,z_{d})^{T}$. Using the independence of $z_{i}-v_{i}$, $z_{i}+v_{i}$, and $z_{j}$, with $j\neq i$, we write the above expectation as
$2\sigma^{2}\bar{S}_{i}={\rm E}\,\left[\left(\frac{\partial f(\boldsymbol{x}_{i})}{\partial z_{i}}\right)^{2}\right]\times{\rm E}\,[(z_{i}-v_{i})^{2}]=2{\rm E}\,\left[\left(\frac{\partial f(\boldsymbol{x}_{i})}{\partial z_{i}}\right)^{2}\right].$
Finally, we observe,
$\gamma_{i}(d)$ $=\int{((f(\boldsymbol{z})-f(\boldsymbol{v}_{\{i\}}{:}\boldsymbol{z}_{\{-i\}}))/(z_{i}-v_{i}))^{2}d\boldsymbol{F}(\boldsymbol{z})dF_{i}(v_{i})}$
$=\int{\left(\frac{\partial f(\boldsymbol{x}_{i})}{\partial z_{i}}\right)^{2}d\boldsymbol{F}(\boldsymbol{z})dF_{i}(v_{i})}$
$=\sigma^{2}\bar{S}_{i}.$
Remark 3.18.
For the activity scores, the relationship between the upper Sobol’ index and the activity score for quadratic functions is an inequality. Indeed, under the same hypothesis of Theorem 3.16, Liu and Owen [13] shows that
$\bar{S}_{i}\leq\frac{\alpha_{i}(d)}{\sigma^{2}},$
for $i=1,\ldots,d$.
4 Numerical Results
In this section we compare four sensitivity measures; Sobol’ indices, DGSMs, activity scores, and global activity scores, numerically, when they are used to conduct the global sensitivity analysis in some examples.
We use the algorithm in Sobol’ [20] to estimate the upper Sobol’ indices, and the Correlation-2 algorithm in Owen [16] for the lower Sobol’ indices. To reduce computational cost, we use the same sample to estimate all indices rather than generating an independent sample for each one. We use this approach for all the sensitivity measures. With this approach, when $N$ Monte Carlo samples are generated, the number of function evaluations is $(d+1)N$ for upper Sobol’, and $(2d+1)N$ for lower Sobol’ indices.
To compute the DGSM (Eqn. (3)) using Monte Carlo, we generate $N$ samples, and for each sample we estimate derivatives using forward finite differences with an increment of $h=0.001$. The total number of function evaluations is $(d+1)N$.
To compute the activity scores $\alpha_{i}(m)$ (Eqn. (9)), we use the same approach as in Constantine et al. [4], and compute the SVD of the matrix $\nabla f(\boldsymbol{x}^{(1)})\cdots\nabla f(\boldsymbol{x}^{(N)})]$, where $x^{(i)},i=1,\ldots,N$, are the Monte Carlo samples, rather than computing the eigenvalue decomposition of $\boldsymbol{C}_{\text{as}}$. We use forward finite differences with an increment of $h=0.001$ to estimate the derivatives. The total number of function evaluations is $(d+1)N$.
To compute the global activity scores $\gamma_{i}(m)$ (Eqn. (13)), we use the same SVD approach as for activity scores. After splitting the inner integral at its singularity, each of the resulting two integrals is approximated using the Gauss-Legendre rule with 10 nodes. The outer integral is estimated with a Monte Carlo sample of size $N$. The total number of function evaluations is $(20d+1)N$.
In the numerical results that follow, we use a Monte Carlo sample size of $N=200,000$ for the Sobol’ indices, DGSM, and activity scores. For the global activity scores, we use a Monte Carlo sample size of $N=10,000$. Computer codes for some of the examples are available at: https://github.com/RuilongYue/global-activity-scores.
An important step in the calculation of activity and global activity scores is determining the dimension $m$ of the active subspace and global active subspace that approximates the original input domain. Consider the eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{d}\geq 0$ of $\Lambda_{\text{as}}$ and $\Lambda_{\text{gas}}$. Define the $m$th normalized cumulative sum of eigenvalues as $\sum_{k=1}^{m}\hat{\lambda}_{k}/\sum_{k=1}^{d}\hat{\lambda}_{k}$, where $\hat{\lambda}_{k}$ is the $k$th estimated eigenvalue. In the following numerical results, for each method, we will pick $m$ as the smallest integer for which the cumulative normalized sum of eigenvalues for the corresponding method is greater than $90\%$.
4.1 Example 1: A test function with noise
Consider the following function
(21) $f(\boldsymbol{z})=\sum_{i=1}^{10}iz_{i}+10(z_{1}z_{2}-z_{9}z_{10})+k\epsilon,\epsilon\sim N(0,1),$
where $\boldsymbol{z}=(z_{1},z_{2},\ldots,z_{10})$. The function has a linear component with inputs in increasing order of importance, and a component with second order interactions. The inputs are independent and have the uniform distribution on (0.5,0.5)(-0.5,0.5). We assume the function is evaluated with some noise modeled by the term $k\epsilon$, where $k$ is a positive constant and $\epsilon$ is a standard normal random variable. Akin to the global sensitivity analysis of stochastic codes (see Fort et al. [6], Hart et al. [7], and Nanty et al. [14]), we want to investigate the accuracy of various global sensitivity measures as the level of noise, $k$, increases. In the numerical results that follow, we consider three cases: no noise ($k=0$), low noise ($k=0.01$), and high noise ($k=1$).
4.1.1 No noise ($k=0$)
Fig. 1 plots the normalized cumulative sum of eigenvalues of $\boldsymbol{C}_{\text{as}}$ and $\boldsymbol{C}_{\text{gas}}$, and the components of the eigenvector that corresponds to the largest eigenvalue for each method. The methods give virtually identical results.
Figure 1: Normalized cumulative sum of eigenvalues (left) and the first eigenvector (right) of $\boldsymbol{C}_{\text{as}}$ and $\boldsymbol{C}_{\text{gas}}$, when $k=0$, in Example 1
Fig. 2 plots Sobol’ lower and upper indices, normalized DGSMs, normalized activity scores, and normalized global activity scores. Normalization is done by dividing each score by the sum of the corresponding scores. We do not normalize Sobol’ indices, for we want to see the unscaled differences between lower and upper Sobol’ indices. For the global activity scores plot, we include the results from $m=1$ and $m=10$. For the activity scores, we only show the results for $m=1$, since $m=10$ for activity scores corresponds to DGSM.
The upper Sobol’ indices and DGSM follow the same pattern. When $m=1$, the activity scores and global activity scores are virtually identical, and they devalue the importance of the first two parameters compared to Sobol’ indices and DGSMs, leading to a discrepancy in the ranking of the least three important inputs. When $m=10$, however, all the methods give the same ranking for the importance of inputs.
(a) Sobol’ indices
(b) Normalized DGSMs
(c) Normalized activity scores
(d) Normalized global activity scores
Figure 2: Sensitivity indices when $k=0$, in Example 1
4.1.2 Low noise ($k=0.01$)
Fig. 3 plots the normalized cumulative sum of eigenvalues of $\boldsymbol{C}_{\text{as}}$ and $\boldsymbol{C}_{\text{gas}}$ in the presence of low noise. We observe the methods have significantly different eigenvalues.
Figure 3: Normalized cumulative sum of eigenvalues (left) and the first eigenvector (right) of $\boldsymbol{C}_{\text{as}}$ and $\boldsymbol{C}_{\text{gas}}$, when $k=0.01$, in Example 1
Fig. 4 plots the sensitivity indices. Based on the cumulative normalized eigenvalue criterion discussed earlier, we pick $m=8$ for the activity scores, and $m=1$ for the global activity scores. With the introduction of noise, we observe that DGSM and activity scores do not look similar to Sobol’ indices (or global activity scores) anymore. Even though DGSM still ranks most of the inputs as the upper Sobol’ indices, the relative differences between the importance of the inputs are very different than that of Sobol’ indices. The activity scores with $m=8$ rank several inputs differently than the Sobol’ upper indices.
The global activity scores, on the other hand, do not seem to be adversely affected by noise. When $m=10$, the global activity scores and the upper Sobol’ indices are virtually indistinguishable. When $m=1$, the global activity scores rank the first three inputs (which are the least important) differently compared to upper Sobol’ indices, and interestingly follow the same pattern as lower Sobol’ indices.
(a) Sobol’ indices
(b) Normalized DGSMs
(c) Normalized activity scores
(d) Normalized global activity scores
Figure 4: Sensitivity indices when $k=0.01$, in Example 1
4.1.3 High noise ($k=1$)
Fig. 5 plots the eigenvalues and the first eigenvector, and Fig. 6 plots the sensitivity indices for the case of high noise. Based on the eigenvalues, we pick $m=8$ for activity scores, and $m=7$ for global activity scores. DGSM and activity scores fail in identifying the important variables. All the inputs have about the same importance according to DGSM, and the least important input becomes the most important according to activity scores. Global activity scores with $m=7$ ranks input parameters the same way as upper Sobol’ indices, except for the ranking of inputs 2 and 4. If we increase $m$ to 10, then global activity scores and the upper Sobol’ indices follow the same pattern and give the same conclusions.
Figure 5: Normalized cumulative sum of eigenvalues (left) and the first eigenvector (right) of $\boldsymbol{C}_{\text{as}}$ and $\boldsymbol{C}_{\text{gas}}$, when $k=1$, in Example 1
(a) Sobol’ indices
(b) Normalized DGSMs
(c) Normalized activity scores
(d) Normalized global activity scores
Figure 6: Sensitivity indices when $k=1$, in Example 1
4.2 Example 2: IEEE 14-bus power system
In this example we consider the global sensitivity analysis of the IEEE 14-bus power system [1]. A description of the problem is provided in Yue et al. [21], who used Sobol’ sensitivity indices, DGSMs, and activity scores to analyze the model. Here we will apply the global activity score method to the power system problem and compare it with the other sensitivity measures.
The output of interest in the sensitivity analysis is the steady state solution for the voltage at busbar 9. The inputs are the active power at each node $i$. We set up the differential equations as in Yue et al. [21], and then find the output by computing the load flow solution with Newton-Raphson method (computational details can be found in Section 6.10 of Saadat [17]). We assume the $i$th input ($i=1,\ldots,14$) follows the distribution $N\left({P}_{i},(0.5{P}_{i})^{2}\right)$, which is one of the uncertainty scenarios considered in Ni et al. [15]. The value of ${P}_{i}$’s are presented in Table 1.
Table 1: Means of normal distributions for power system
Busbar i{i} 1 2 3 4 5 6 7
${P}_{i}$ 0 21.7 94.2 47.8 7.6 11.2 0
Busbar i{i} 8 9 10 11 12 13 14
${P}_{i}$ 0 29.5 9 3.5 6.1 13.5 14.9
Figure 7 plots the sensitivity indices. DGSM and activity score methods suggest very different important variables than Sobol’ indices. The normalized DGSM and activity scores for nodes 4, 6, 13, and 14 are about zero (the largest value is $3\times 10^{-5}$). However, these nodes have significant importance, according to Sobol’ indices. For example, node 4 has the second largest upper Sobol’ index with a value of 0.22. The upper Sobol’ indices for nodes 6 and 14 are 0.09 and 0.082. The global activity scores, on the other hand, are very similar to upper Sobol’ indices. According to upper Sobol’ indices, the six most important inputs, in decreasing order, are 3, 4, 9, 13, 6, 14, and according to global activity scores ($m=10$), the most important inputs are 3, 9, 4, 13, 14, 6.
(a) Sobol’ indices
(b) Normalized DGSMs
(c) Normalized activity scores
(d) Normalized global activity scores
Figure 7: Sensitivity indices for power system
4.3 Example 3: A counterexample from Sobol’ and Kucherenko
Sobol’ and Kucherenko [10] present a function where the ranking of important variables by DGSM may result in false conclusions and give a different ranking than the upper Sobol’ indices. The function is given by
$f(\boldsymbol{x})=\sum_{i=1}^{4}c_{i}\left(x_{i}-1/2\right)+c_{12}\left(x_{1}-1/2\right)\left(x_{2}-1/2\right)^{5},$
where $\boldsymbol{x}=(x_{1},x_{2},x_{3},x_{4}),c_{i}=1,1\leq i\leq 4$ and $c_{12}=50$. The upper Sobol’ index for $x_{1}$ and $x_{2}$ is 0.289, and the upper Sobol’ index for $x_{3}$ and $x_{4}$ is 0.237.
We want to examine how the global activity scores compare with upper Sobol’ and DGSM for this example. Fig. 8 plots the sensitivity indices for the function. Normalized DGSMs and activity scores ($m=2$) give the same results, and they incorrectly find the first input variable less important than the second. Sobol’ and Kucherenko explain the reason for this discrepancy by the strong nonlinearity of the term $c_{12}\left(x_{1}-1/2\right)\left(x_{2}-1/2\right)^{5}$. The normalized global activity scores of the input variables are much closer to the Sobol’ sensitivity indices, even though the global activity score of the first variable is slightly larger than the second one.
(a) Sobol’ indices
(b) Normalized DGSMs
(c) Normalized activity scores
(d) Normalized global activity scores
Figure 8: Sensitivity indices for the example from Sobol’ and Kucherenko
5 Conclusions
Through the numerical results, we observed that the global activity scores align with the upper Sobol’ indices much more closely than DGSM or activity scores in the presence of noise (Example 1), large variance (Example 2), and strong nonlinearity (Example 3). In the absence of such factors, we found empirically that the global activity scores behave similarly to DGSM and activity scores (we report some of these results - specifically the no-noise case in Example 1 - in the paper). These findings highlight the robustness of global activity scores in challenging settings where other sensitivity measures may mislead. Beyond the examples presented here, we expect the method to be broadly applicable in problems involving noisy simulations or strong nonlinearities, making it a valuable addition to the global sensitivity analysis toolbox.